OpenStudy (andpeggy):

a drugist has two solutions of alcohol. one is 25% alcohol. the other is 45% alcohol. he wants to mix these two solutions to get 36 ounces that will be 30% alcohol. how many ounces of each of these two solutions should he mix together?

OpenStudy (wolf1728):

Set up two equations A) x + y = 36 B) .25x + .45y = (.30 * 36) Then solve for x and y

OpenStudy (wolf1728):

Do you know how to solve for x and y?

OpenStudy (andpeggy):

not really, are you supposed to multiply .25 to the first equation?

OpenStudy (wolf1728):

That would be one way but I would multiply equation "B" by minus 4

OpenStudy (andpeggy):

where did the minus 4 come from?

OpenStudy (wolf1728):

We'd want to have it so that one of the unknowns could be eliminated. Multiplying equation "B" by -4 would make it B) -x -1.8Y = -43.20 When we add that to equation "A" the "x" term gets eliminated

OpenStudy (wolf1728):

Do you want to see how that's done?

OpenStudy (andpeggy):

yes please

OpenStudy (wolf1728):

A) x + y = 36 B) -x -1.8 y = -43.20 Adding both equations we get -.80y = -7.2

OpenStudy (andpeggy):

ohhh I see I kept multiplying the minus 4 to the a equation as well

OpenStudy (andpeggy):

so y would turn out to equal 9?

OpenStudy (wolf1728):

No you multiply equation B) .25x + .45y = (.30 * 36) by -4 and it becomes B) -x -1.8 y = -43.20 THen add it to equationA And yes "y" is nine!!!

OpenStudy (andpeggy):

and so to get x would you just do 36-9?

OpenStudy (andpeggy):

which is 27 but is that how id have to solve x?

OpenStudy (wolf1728):

That is exactly how it is done

OpenStudy (wolf1728):

So basically you need 27 ounces of 25% and 9 ounces of 45%

OpenStudy (andpeggy):

ok thank you so so much for helping me out! <3 :D

OpenStudy (wolf1728):

u r welcome